BINOMIAL COEFFICIENTS (20 pts)

1.    Expand the (2𝑚 − 2𝑏)3 using binomial coefficient. (14 pts)

2.    Find for the coefficient of a5b5; (a - 4b )10 (3 pts)

3.    Find the 5th term after expanding the expression (3x – 4y)15 (3 pts)

 

PIGEONHOLE PRINCIPLE

Show that in a group of 27 English words, there must be at least two that begin with the same letter. (5 pts) 

1
Expert's answer
2021-05-17T16:53:31-0400

Solution.

1.

"(2x-2y)^3=8x^3-3•4x^2•2y+3•2x•4y^2-8y^3=\newline 8x^3-24x^2y+24xy^2-8y^3."

2.

"(a-4b)^{10}=a^{10}-10a^9\u20224b+45a^8\u2022(4b)^2-120a^7\u2022(4b)^3+210a^6\u2022(4b)^4-252a^5\u2022(4b)^5+210a^4\u2022(4b)^6-120a^3\u2022(4b)^7+45a^2\u2022(4b)^8-10a\u2022(4b)^9+(4b)^{10}\n."

Answer.The coefficient of "a^5b^5:" -258048.

3.

"(3x-4y)^{15}."

The 5th term after expanding: "\u0422_5=\u0421_4^{15}(3x)^{11}(4y)^{4}=1365\u2022177147x^{11}\u2022256y^4=1365\u2022177147\u2022256x^{11}y^4."

PIGEONHOLE PRINCIPLE

The pigeonhole principle states that if n items are put into m containers, with n>m, then at least one container must contain more than one item. We have n=27 English words and m=26 English letters. n>m,27>26, there are at least two words that begin with the same letters.